OpenStudy (anonymous):

Determine absolute minimum value of f(x)=2+3sin(x)cos(x) on the interval [0,pi]

OpenStudy (anonymous):

f(x)=2+3sin(x)cos(x) f(x)=2+(3/2)sin(2x) min value of sin(2x) =-1 hence min value of f(x) is 2-3/2 =1/2